Stoichiometry

How many grams would 8.1  1021 molecules of sucrose (C12H22O11) weigh?
1
Expert's answer
2021-09-26T12:14:48-0400

Number of molecules of sucrose = 8.1 ×10²¹ molecules

Mass of sucrose = ?

Solution:

First of all we will calculate the number of moles by using Avogadro number.

1 mole × 8.1 ×10²¹ molecules / 6.022×10²³ molecules

1.35 × 10⁻² mol

Mass of sucrose:

Mass = number of moles × molar mass

Molar mass = 342.3 g/mol

Mass = 1.35 × 10⁻² mol ×342.3 g/mol

Mass = 462.1 × 10⁻² g

Mass = 4.6 g